# Superposition Prompting: Improving and Accelerating Retrieval-Augmented Generation

### v1

@@ 4.1.1 NaturalQuestions-Open
We present speedup vs. accuracy comparisons in Table 1. For the TF-IDF baseline, we use TF-IDF (from SciPy package Virtanen et al., 2020) to select the top-$k$ documents conditioned on the query, then perform “naive LLM-RAG” (as described in Section 3.1). Our BM-25 baseline is equivalent, except we use Brown, 2020 for the top-$k$ document selection. We also have an equivalent baseline where we use Contriever (Izacard et al., 2021) to select the top-$k$ documents. 666 For a more generous representation of the BM-25, TF-IDF, and Contriever baselines, we compute the speedup metrics assuming document KV caching (although to our knowledge, this has not been previously proposed in literature). Note that caching is not possible with Naive LLM-RAG or Attention Sort since the order of documents is variable, and in general, documents attend to other documents (thus also parallelization is not possible). We compare against the recently proposed Attention Sort method, using their method exactly as described in Peysakhovich & Lerer, 2023. Finally, we compare against Prompt Cache (Gim et al., 2023). Note that Naive LLM-RAG, Prompt Cache, and Attention Sort always attend to all documents.
In addition to Table 1, we present various architectural ablation studies in Appendix A to justify our design decisions.
@@ 4.1.2 MuSiQue
A slight modification is made to superposition prompting to handle the multi-hop reasoning setup of MuSiQue. Specifically, we iteratively apply superposition pruning to build a chain of $t\times k$ documents777 Note that only the first $k$ documents chosen are cacheable. Subsequent documents are not cacheable since their KV caches depend on preceding documents (which are dynamically chosen during query serving). , where $t$ and $k$ are hyperparameters. At each time step $\{1,\ldots t\}$, we create a superposition with all remaining documents, prune to keep the top $k$, prepend those (cached) documents to the running prefix, then repeat. A visual depiction of this iterative superposition is presented in Figure 7. We hypothesize that iterative superposition can improve performance since we equip the LLM to iteratively solve the multi-hop reasoning challenge.
@@ 4.2.2 Superposition Prompting Can Improve Accuracy
One explanation for the accuracy improvement is how superposition prompting reduces sequence lengths as perceived by the transformer. Recent studies have investigated the apparent lack of “length extrapolation” abilities of transformer-based LLMs (Press et al., 2021; Ruoss et al., 2023; Kazemnejad et al., 2023). One convenient property of superposition prompting is that—from the perspective of the transformer—the maximum sequence length observed is the longest path through the graph.888 This means the effective (perceived) sequence length is $\mathcal{O}(1)$ instead of $\mathcal{O}(n_{d})$, where $n_{d}$ is the number of offline documents. For example, with NaturalQuestions-Open, superposition prompting decreases the maximum path (and thus the sequence length) from an average of 2923 tokens to 206 tokens. In this sense, superposition prompting for RAG can enable non-long-context transformers to perform well on long sequences. This property could allow model developers to significantly reduce pretraining costs (since training special-purpose “long-context” LLMs leads to increased costs (Press et al., 2021)).
Another explanation for the accuracy improvement is the LLM “distraction” phenomenon. The previous works of Liu et al., 2023a; Borgeaud et al., 2021a; Shi et al., 2023 present arguments for how LLMs can be sensitive to noisy or irrelevant context. With the inclusion of the path pruning mechanism, we equip the model with a structured way to filter out such noise.
@@ 5 Ablations
We present ablations against our design decisions here as well as extensively in Appendix A.
@@ 5.1 Position Assignment Ablation
In Table 3, we investigate the effect of the position assignment strategy during superposition prompting. We compare our proposed equilibrium path positioning to the left aligned strategy described in Section 3.2.2. Our findings validate our hypothesis outlined in Section B.2, where we speculated that left alignment would result in worse performance (due to long sequence attention bias).
@@ 5.2 Path Saliency Metric Ablation
In Table 4, we ablate our choice of “path saliency” metric. We compare against two other baselines—attention and none. With none, we simply do not prune. The attention baseline consists of using the attention scores for each document (average across tokens, layers and attention heads) as the path score. We highlight that our Bayesian path saliency significantly outperforms attention-based scoring.
@@ Acknowledgements
Appendix A Additional Ablations
A.1 Superposition Factor Ablation
We introduce the hyperparameter superposition factor as a way to interpolate between a fully superimposed and fully “classical” prompt. Larger superposition factors correspond to “more superimposed” prompts, whereas smaller superposition factors correspond to “less superimposed” prompts (achieved by combining adjacent documents before creating prompt paths).
Formally, define $m$ as the number of documents considered for a retrieval-augmented generation query (for instance, this is $m=20$ for common settings of NaturalQuestions-Open (Liu et al., 2023a) and MuSiQue (Trivedi et al., 2022)). By setting a superposition factor $\gamma\in[1,m]$, we compute the “effective documents per path” as $\lfloor\frac{m}{\gamma}\rceil$. Importantly, note that when $\gamma=1$, we’ve reduced to the “classical” (Naive LLM-RAG) case. We perform an ablation by sweeping this superposition factor parameter and present results in Table 5.
Figure 5: Visual depiction of the superposition factor, which is measure of how “superpositioned” a prompt is. Refer to Section A.1 for a definition.
Figure 6: Sweeping values of superposition factor on the NaturalQuestions-Open dataset with a variety of models (sourced from Table 5).
Table 5: Retrieval augmented generation accuracy for various models on the NaturalQuestions-Open dataset varying the superposition factor. We observe that, in general, greater superposition results in better performance vs. accuracy tradeoff.
Compute Cycles Theor. Speedup Accuracy
Model Superposition Factor
bloomz-3b 1 9.21e+12 1.1 0.005
2 4.07e+12 2.4 0.191
4 1.79e+12 5.5 0.210
10 5.09e+11 19.1 0.227
20 9.92e+10 98.2 0.223
bloomz-7b1 1 2.10e+13 1.1 0.022
2 9.50e+12 2.3 0.269
4 4.23e+12 5.3 0.260
10 1.22e+12 18.2 0.241
20 2.38e+11 93.4 0.253
mpt-7b-instruct 1 2.02e+13 1.1 0.026
2 9.18e+12 2.4 0.398
4 4.09e+12 5.3 0.389
10 1.18e+12 18.4 0.382
20 2.31e+11 94.0 0.465
A.2 Top-$k$ Ablation
In Table 6, we sweep values for top-$k$ for our method, where $k$ are the number of documents retained for generating the answer. For completeness, we also ablate this hyperparameter for all “ranking-based” baselines, namely BM-25, TF-IDF, and Contriever (described in Section 4.1.1). Note that for superposition prompting, this top-$k$ value corresponds the number of paths retained after path pruning. We see that accuracy tends to peak around $k=2$ to $k=4$, although it comes at a cost to speedup versus $k=1$. Interestingly we do see that performance decreases steadily for increasing $k>4$. This seems to coincide with the arguments put forth in (Shi et al., 2023) and Press et al., 2021 and Liu et al., 2023a.
We also report CUDA speedup values for our ablations, using methodology further explained in Appendix C.
Table 6: Retrieval augmented generation accuracy and speedups for various models on the NaturalQuestions-Open dataset, sweeping top-$k$.
Compute Cycles Theor. Speedup CUDA Speedup Accuracy
@@ A.2 Top-$k$ Ablation
A.3 Runtime Optimization Ablation
In Table 7 we present a full breakdown of the incremental effect of the optimizations proposed in Section 3.2.3 and Section 3.3.
Appendix B Supplementary Algorithm Details
B.1 Bayesian Path Selection
Define $H_{\triangleright}:(\mathbb{R}^{*\times n_{v}},\mathcal{S}^{*})\rightarrow\mathbb{R}$ as the “shifted” cross entropy between a logit tensor and sequence tensor of the same sequence dimension (where we discard element 1 from the input sequence and the last element from the logit tensor). Formally, for our ForkJoin prompt topology, we compute the Bayesian path saliency as follows.
Table 7: Ablation of speedup vs. accuracy on the NaturalQuestions-Open dataset by enabling/disabling the optimizations proposed in Section 3.3
@@ B.1 Bayesian Path Selection
Model Path Pruning Path Caching Path Parallel.
@@
As detailed in Section B.3, during superposition prompting inference, we compute logits for the preamble $\pi\in\mathbb{R}^{\|\boldsymbol{p}\|\times n_{v}}$ (where $n_{v}$ is the vocabulary size), logits for all documents $\delta_{i}\in\mathbb{R}^{\|\boldsymbol{d}_{i}\|\times n_{v}}$, and logits for all queries $\phi_{i}\in\mathbb{R}^{\|\boldsymbol{q}\|\times n_{v}}$ for $i\in D$. Then, for each $k\in D$, we can compute
@@
$=\frac{\log H_{\triangleright}(\delta_{i},\boldsymbol{d}_{i})}{\|\boldsymbol{d}_{i}\|}+\frac{\log H_{\triangleright}(\pi,\boldsymbol{p})}{\|\boldsymbol{p}\|}-C$ (2b)
Equation 2a follows from Bayes Rule (where $C$ is some unspecified constant), while Equation 2b follows from our definition of language model. The $C$ term is inconsequential for our use, since we eventually softmax before comparing the likelihoods. This justifies why we duplicate queries in the ForkJoin topology—without duplicating the query for each path, we only have access to $P(\boldsymbol{q}_{i}\mid\boldsymbol{d}_{1},\ldots\boldsymbol{d}_{n_{d}},\boldsymbol{p})$, not the independent terms $\{P(\boldsymbol{q}_{i}\mid\boldsymbol{d}_{i},\boldsymbol{p})\mid i\in D\}$ We choose to normalize the log likelihood terms by their corresponding sequence lengths ($\|\boldsymbol{d}_{i}\|$ and $\|\boldsymbol{q}\|$) to effectively achieve a “likelihood density per token,” which prevents bias against shorter sequence lengths999 Note that SGPT (Muennighoff, 2022) avoided length bias by truncating sequences to a common length. We choose not to follow this design decision to prevent potential loss of vital information in the input data. .
@@ B.2 Equilibrium Position Assignment
Section 3.2.2 outlined the intuition behind the equilibrium position assignment algorithm.
@@
  $K:=\operatorname*{argmin}^{K}_{i\in D}P(\boldsymbol{d}_{i}\mid\boldsymbol{q}_{i},\boldsymbol{p})$triangleright Using Section B.1.
@@
B.3 Full Algorithm Outline
[…]
  • $\boxed{x}\in\mathcal{M}$ (KV cache used as input)
  • $\boldsymbol{y}\in\mathcal{S}$ (new tokens not included in KV cache)
Following the insight of Section 3.3.2, we also define
B.4 Real-Valued Position Assignments
For reasons outlined in Section 3.2.2, superposition prompting is very naturally suited for transformers which accept continuously-valued token position assignments (i.e. the position interpolation property as defined by Chen et al., 2023). The ALiBi (Press et al., 2021) transformer positional encoding scheme supports this requirement naturally (Al-Khateeb et al., 2023) and is used in numerous open source LLM models.
Other popular positional encoding schemes, such as Rotary Position Encoding (Su et al., 2021), have been shown to be able to exhibit this property (Chen et al., 2023).101010 To be clear, we are interested in the interpolation property not for the sake of context length extrapolation as other works are, but for the interpolation itself. In fact, we have been able to apply superposition prompting to the RoPE-based LLaMa 2 7B-Instruct model (Touvron et al., 2023b), where it achieves 0.366 Best EM Subspan on Natural Questions using top-$1$ path pruning (without fine-tuning).
B.5 Iterative Superposition
We present a visual depiction of iterative superposition in Figure 7.
@@ Appendix C CUDA Benchmarks
Table 8: CUDA speedup measurements for remaining baselines methods not enumerated in Table 6. This table is meant to be directly comparable to Table 6.
@@
We notice that the actual measured speedups are an order of magnitude smaller than the theoretical maximum speedups calculated. This is expected—as we heavily optimize the FLOPs (up to 100$\times$), the memory bottlenecks become much more apparent. We expect that a fused CUDA kernel implementation could bridge the order of magnitude gap, similar to how Dao et al., 2022 achieves an order of magnitude improvement over the naive PyTorch implementation by mitigating memory transfer bottlenecks.
Appendix D Dataset Statistics

### v2

@@ 4.1.1 NaturalQuestions-Open
OpenELM-3B-In. Naive LLM-RAG 1.03e+13 1.0 0.001
BM-25 1.07e+11 96.9 0.166
TF-IDF 1.07e+11 96.9 0.215
Contriever 5.62e+11 18.4 0.191
Attention Sort 3.09e+13 0.3 0.000
Prompt Cache 1.25e+11 82.4 0.000
Superposition (Ours) 1.07e+11 96.8 0.241
@@
We present speedup vs. accuracy comparisons in Table 1. For the TF-IDF baseline, we use TF-IDF (from SciPy package Virtanen et al., 2020) to select the top-$k$ documents conditioned on the query, then perform “naive LLM-RAG” (as described in Section 3.1). Our BM-25 baseline is equivalent, except we use Brown, 2020 for the top-$k$ document selection. We also have an equivalent baseline where we use Contriever (Izacard et al., 2021) to select the top-$k$ documents.777 For a more generous representation of the BM-25, TF-IDF, and Contriever baselines, we compute the speedup metrics assuming document KV caching (although to our knowledge, this has not been previously proposed in literature). Note that caching is not possible with Naive LLM-RAG or Attention Sort since the order of documents is variable, and in general, documents attend to other documents (thus also parallelization is not possible). We compare against the recently proposed Attention Sort method, using their method exactly as described in Peysakhovich & Lerer, 2023. Finally, we compare against Prompt Cache (Gim et al., 2023). Note that Naive LLM-RAG, Prompt Cache, and Attention Sort always attend to all documents.
In addition to Table 1, we present various architectural ablation studies in Section 5 and Appendix A to justify our design decisions.
@@ 4.1.2 MuSiQue
A slight modification is made to superposition prompting to handle the multi-hop reasoning setup of MuSiQue. Specifically, we iteratively apply superposition pruning to build a chain of $t\times k$ documents888 Note that only the first $k$ documents chosen are cacheable. Subsequent documents are not cacheable since their KV caches depend on preceding documents (which are dynamically chosen during query serving). , where $t$ and $k$ are hyperparameters. At each time step $\{1,\ldots t\}$, we create a superposition with all remaining documents, prune to keep the top $k$, prepend those (cached) documents to the running prefix, then repeat. A visual depiction of this iterative superposition is presented in Figure 6. We hypothesize that iterative superposition can improve performance since we equip the LLM to iteratively solve the multi-hop reasoning challenge.
@@
OpenELM-3B-In. Naive LLM-RAG 8.81e+12 1.0 0.006 0.000
Attention Sort 2.64e+13 0.3 0.009 0.001
Prompt Cache 1.61e+11 54.7 0.000 0.000
Superposition (X=1, K=4) (Ours) 1.42e+11 62.1 0.028 0.000
Superposition (X=2, K=4) (Ours) 5.97e+11 14.8 0.039 0.006
Superposition (X=4, K=1) (Ours) 1.99e+12 4.4 0.059 0.019
@@ 4.2.2 Superposition Prompting Can Improve Accuracy
One explanation for the accuracy improvement is how superposition prompting reduces sequence lengths as perceived by the transformer. Recent studies have investigated the apparent lack of “length extrapolation” abilities of transformer-based LLMs (Press et al., 2021; Ruoss et al., 2023; Kazemnejad et al., 2023). One convenient property of superposition prompting is that—from the perspective of the transformer—the maximum sequence length observed is the longest path through the graph.999 This means the effective (perceived) sequence length is $\mathcal{O}(1)$ instead of $\mathcal{O}(n_{d})$, where $n_{d}$ is the number of offline documents. For example, with NaturalQuestions-Open, superposition prompting decreases the maximum path (and thus the sequence length) from an average of 2923 tokens to 206 tokens. In this sense, superposition prompting for RAG can enable non-long-context transformers to perform well on long sequences. This property could allow model developers to significantly reduce pretraining costs (since training special-purpose “long-context” LLMs leads to increased costs (Press et al., 2021)).
Another explanation for the accuracy improvement is the LLM “distraction” phenomenon. The previous works of Liu et al., 2023a; Borgeaud et al., 2021a; Shi et al., 2023 present arguments for how LLMs can be sensitive to noisy or irrelevant context. With the inclusion of the path pruning mechanism, we equip the model with a structured way to filter out the “noise” (i.e. irrelevant documents).
4.2.3 Sensitivity to ALiBi vs. RoPE
For reasons outlined in Section 3.2.2, superposition prompting is very naturally suited for transformers which accept continuously-valued token position assignments (i.e. they exhibit the position interpolation property as defined by Chen et al., 2023). While the ALiBi positional encoding scheme has been shown to posses this property, it has been suggested that fine-tuning may be required to equip Rotary Position Embedding (RoPE)-based models with this property.
Our experiments validate that our proposed equilibrium positional assignment mechanism is compatible even with a non-fine-tuned RoPE-based model (i.e. the OpenELM family). We leave it to future studies to measure the extent to which fine-tuning may improve accuracy (if at all).
We note that OpenELM-3B-Instruct has significantly lower accuracy for many baselines, such as AttentionSort, Naive LLM-RAG and even Prompt Cache. We hypothesize that this is due to the lack of length extrapolation capabilities of RoPE, which would become more pronounced for those baselines.
@@ 5.1 Position Assignment Ablation
In Table 3, we investigate the effect of the position assignment strategy during superposition prompting. We compare our proposed equilibrium path positioning to the left aligned strategy described in Section 3.2.2. Our findings validate our hypothesis outlined in Algorithm 1, where we speculated that left alignment would result in worse performance (due to long sequence attention bias).
@@ 5.2 Path Saliency Metric Ablation
In Table 4, we ablate our choice of “path saliency” metric. We compare against two other baselines—attention and none. With none, we simply do not prune. The attention baseline consists of using the attention scores for each document (average across tokens, layers and attention heads) as the path score. We highlight that our Bayesian path saliency significantly outperforms attention-based scoring, as well as the control baseline.
@@
5.3 Superposition Factor Ablation
We introduce the hyperparameter superposition factor as a parameter to interpolate between a fully superimposed and fully “classical” prompt. Larger superposition factors correspond to “more superimposed” prompts, whereas smaller superposition factors correspond to “less superimposed” prompts (achieved by combining adjacent documents before creating prompt paths).
Formally, we define $m$ as the number of documents considered for a retrieval-augmented generation query (for instance, this is $m=20$ for common settings of NaturalQuestions-Open (Liu et al., 2023a) and MuSiQue (Trivedi et al., 2022)). By setting a superposition factor $\gamma\in[1,m]$, we compute the “effective documents per path” as $\lfloor\frac{m}{\gamma}\rceil$. Importantly, note that when $\gamma=1$, we’ve reduced to the “classical” (Naive LLM-RAG) case. We perform an ablation by sweeping this superposition factor parameter and present results in Figure 5. A visual representation is presented in Figure 7.
The curves generally show improvements along both axes as we increase the superposition quotient. Interestingly, the maximal accuracy may not be fully superimposed, suggesting that this value should be tuned for the given application.
Figure 5: Sweeping values of superposition factor (SF) on the NaturalQuestions-Open dataset with a variety of models.
@@ Acknowledgements
Appendix A Ablations (cont.)
A.1 Top-$k$ Ablation
Here, we sweep values for top-$k$ for our method, where $k$ are the number of documents retained for generating the answer (full table results are provided in Table 5). For completeness, we also ablate this hyperparameter for all “ranking-based” baselines, namely BM-25, TF-IDF, and Contriever (described in Section 4.1.1). Note that for superposition prompting, this top-$k$ value corresponds the number of paths retained after path pruning. We see that accuracy tends to peak around $k=2$ to $k=4$, although it comes at a cost to speedup versus $k=1$. Interestingly we do see that performance decreases steadily for increasing $k>4$. This seems to coincide with the arguments put forth in (Shi et al., 2023) and Press et al., 2021 and Liu et al., 2023a of “increased distraction” as the context length increases.
Table 5: Retrieval augmented generation accuracy and speedups for various models on the NaturalQuestions-Open dataset, sweeping top-$k$. Note that “Comp.” is used as an abbreviation for “Compute”, and “Sp.” is used as an abbreviation for “Speedup”.
Comp. Cycles Theor. Sp. CUDA Sp. Accuracy
@@ A.1 Top-$k$ Ablation
A.2 Runtime Optimization Ablation
In Table 6 we present a full breakdown of the incremental effect of the path pruning, path caching, and path parallelism optimizations proposed in Section 3.2.3, Section 3.3.1, and Section 3.3.2, respectively.
Table 6: Ablation of speedup vs. accuracy on the NaturalQuestions-Open dataset by enabling/disabling the optimizations proposed in Section 3.3. Each of Pruning?/Caching?/Parallelism? correspond to path pruning, path caching, and path parallelism enabled.
@@ A.2 Runtime Optimization Ablation
Model Pruning? Caching? Parallelism?
OpenELM-3B-In. False False False 9.81e+12 1.1 0.005
True 6.82e+11 15.1 0.005
True False 9.86e+11 10.5 0.005
True 1.18e+11 87.2 0.005
True False False 9.80e+12 1.1 0.241
True 6.71e+11 15.4 0.241
True False 9.74e+11 10.6 0.241
True 1.07e+11 96.9 0.241
@@
Appendix B Supplementary Algorithm Details
B.1 Bayesian Path Selection
Define $H:(\mathbb{R}^{*\times n_{v}},\mathcal{S}^{*})\rightarrow\mathbb{R}$ as the language modeling cross-entropy. 101010 More specifically, this is the “shifted” cross entropy between a logit tensor and sequence tensor of the same sequence dimension (where we discard element 1 from the input sequence and the last element from the logit tensor). Formally, for our ForkJoin prompt topology, we compute the Bayesian path saliency as follows.
As detailed in Section B.3, during superposition prompting inference, we compute logits for the preamble $\pi\in\mathbb{R}^{\|\boldsymbol{p}\|\times n_{v}}$ (where $n_{v}$ is the vocabulary size), logits for all documents $\delta_{i}\in\mathbb{R}^{\|\boldsymbol{d}_{i}\|\times n_{v}}$, and logits for all queries $\phi_{i}\in\mathbb{R}^{\|\boldsymbol{q}\|\times n_{v}}$ for $i\in D$. Then, for each $k\in D$, we can compute:
@@ B.1 Bayesian Path Selection
$=\frac{\log H(\delta_{i},\boldsymbol{d}_{i})}{\|\boldsymbol{d}_{i}\|}+\frac{\log H(\pi,\boldsymbol{p})}{\|\boldsymbol{p}\|}-C$ (2b)
Equation 2a follows from Bayes Rule (where $C$ is some unspecified constant), while Equation 2b follows from our definition of language model. The $C$ term is inconsequential for our use, since we eventually softmax before comparing the likelihoods. This justifies why we duplicate queries in the ForkJoin topology—without duplicating the query for each path, we only have access to $P(\boldsymbol{q}_{i}\mid\boldsymbol{d}_{1},\ldots\boldsymbol{d}_{n_{d}},\boldsymbol{p})$, not the independent terms $\{P(\boldsymbol{q}_{i}\mid\boldsymbol{d}_{i},\boldsymbol{p})\mid i\in D\}$ We choose to normalize the log likelihood terms by their corresponding sequence lengths ($\|\boldsymbol{d}_{i}\|$ and $\|\boldsymbol{q}\|$) to effectively achieve a “likelihood density per token,” which prevents bias against shorter sequence lengths111111 Note that SGPT (Muennighoff, 2022) avoided length bias by truncating sequences to a common length. We choose not to follow this design decision to prevent potential loss of vital information in the input data. .
@@ B.2 Equilibrium Position Assignment
Section 3.2.2 outlined the intuition behind the equilibrium position assignment algorithm. Here, we algorithmically describe the method.
@@
B.3 Full Algorithm Outline
[…]
  • $\boxed{x}\in\mathcal{M}$ (KV cache used as input)
  • $\boldsymbol{y}\in\mathcal{S}$ (new tokens not included in KV cache)
Following the insight of Section 3.3.2, we also define
@@ B.3 Full Algorithm Outline
  $K:=\operatorname*{argmin}^{K}_{i\in D}P(\boldsymbol{d}_{i}\mid\boldsymbol{q}_{i},\boldsymbol{p})$triangleright Using Equation 2.
@@ Appendix C CUDA Benchmarks
Table 7: CUDA speedup measurements for remaining baselines methods not enumerated in Table 5 (this table is meant to be directly comparable to Table 5).
@@
We notice that the actual measured speedups are an order of magnitude smaller than the theoretical maximum speedups calculated. This is expected—as we heavily optimize the FLOPs (up to 100$\times$), the memory bottlenecks begin to dominate the runtime. We expect that a fused CUDA kernel implementation could bridge the order of magnitude gap, similar to how Dao et al., 2022 achieves an order of magnitude improvement over the naive PyTorch implementation by mitigating memory transfer bottlenecks.
Appendix D Additional Visualizations
D.1 Iterative Superposition
We present a visual depiction of iterative superposition (from Section 4.1.2) in Figure 6.
D.2 Superposition Factor
We present a visual depiction of superposition factor (from Section 5.3) in Figure 7.
Figure 7: Visual depiction of the superposition factor, which is measure of how “superpositioned” a prompt is. A rigorous definition is provided in Section 5.3.
Appendix E Dataset Statistics